Circular No. 8736

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006du, 2006dx, 2006dy, 2006dz, 2006ea, 2006eb, 2006ec

Six additional apparent supernovae have been discovered on CCD images (all unfiltered except for 2006dz): 2006dx by the “Nearby Supernova Factory” collaboration (cf. CBET 584) on NEAT images obtained at Palomar; 2006dy by K. Itagaki (communicated by S. Nakano; cf. IAUC 8700); 2006dz by C. Contreras and G. Folatelli (Carnegie Supernova Project) on images obtained in six different optical bands at Las Campanas Observatory; 2006ea and 2006ec by N. Joubert, D. Madison, D. Winslow, and W. Li (LOSS/KAIT; cf. IAUC 8733); and 2006eb by R. Quimby (cf. IAUC 8622, 8723).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006dx July 19.34 21^{h}10^{m}34.^{s}49 −21^{o}34′42′′.7 19.2  –
2006dy July 25.54 14 22 11.45 +13 55 14.2 16.6 10′′ E, 9′′ N
2006dz May 31.14 13 30 00.37 +13 24 43.1 20.0 30′′ W, 15′′ S
2006ea July 25.35 22 39 37.91 +11 46 16.0 18.8 6′′.2 E, 3′′.9 N
2006eb July 22.41 1 13 41.10 − 0 06 06.4 17.5 10′′.5 W, 3′′.8 N
2006ec Aug. 4.43 23 38 12.46 −20 46 48.6 16.7 10′′.5 W, 7′′.0 N

SNe 2006du (cf. IAUC 8733) and 2006dx are both type-II supernovae (details given on CBET 584). Additional available magnitudes for 2006dy in NGC 5587: 2002 Jan. 19, [19.0 (Itagaki); 2006 June 28, [19.5 (Itagaki); July 22.22 UT, 18.5 (Li, via KAIT); 26.20, 16.5 (Li, via KAIT). Itagaki adds that nothing is visible at the position of 2006dy on the Digitized Sky Survey. Li reports the following position end figures for 2006dy from a KAIT image taken on July 26.20: 11^{s}.55, 13′′.8. SN 2006dy is a type-Ia supernova, some 10 days before maximum light on July 26 (cf. CBET 587). Further to the tabulated Sloan r′ magnitude above, additional magnitudes for 2006dz in NGC 5185: May 17, [21.7; June 4.10, 19.0; 5.15, 19.0; 7.10, 18.9; 8.09, 18.9; 11.07, 19.0; 12.09, 19.0 (additional details given on CBET 588). Additional KAIT magnitudes for 2006ea in UGC 12134: 2005 Nov. 20.12, [19.5; 2006 July 10.36, [18.7; 26.39, 18.8. SN 2006ea is probably a type-Ib event, ∼ 1 week past maximum on July 27 (details on CBET 589). Additional magnitudes for 2006eb in UGC 771: 2005 Aug. 22-Dec. 3 (co-addition of images), [19.0; 2006 July 24.41, 17.5. SN 2006eb is a type-Ia supernova some weeks past maximum (details on CBET 592). Additional KAIT magnitudes for 2006ec in MCG –04-55-14: July 3.44, [18.6; Aug. 5.46, 17.0. A KAIT image from July 16.45 also shows a hint of 2006ec (limiting mag 17.4).

2006 August 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green